OpenStudy (anonymous):

Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

zepdrix (zepdrix):

Oh I see this same question was asked somewhere else already. Were you able to figure it out? :)

OpenStudy (anonymous):

Take Summation of Forces Horizontal and vertical EFX = 200cos60 - 100cos(180-170) = 1.519224699 EFY = 200sin60 100sin(180-170) = 190.5678785 Resultant/Magnitude R2 = (ΣFx)² (ΣFy)² R = 190.5759541 THETA¸ =arctan [EFY2 /EFX2] THETA¸ =89.54324728

OpenStudy (anonymous):

is tht right?

zepdrix (zepdrix):

Mmm so sloppy :) Lemme work it out real quick heh. \[\Large \vec{F_1} \quad=\quad \left<\color{orangered}{200\cos60}, \color{royalblue}{200\sin60}\right>\]\[\Large \vec{F_2} \quad=\quad \left<\color{orangered}{100\cos170}, \quad\color{royalblue}{100\sin170}\right>\] \[\Large \vec F \quad=\quad \vec F_1+ \vec F_2\]\[\large \vec F \quad=\quad \left<\color{orangered}{200\cos60+100\cos170},\quad\color{royalblue}{200\sin60+100\sin170}\right>\] Which simplifies to,

zepdrix (zepdrix):

Mmm ya looks like you're on the right track so far!\[\large \vec F \quad=\quad \left<\color{orangered}{1.519},\quad\color{royalblue}{190.57}\right>\]

zepdrix (zepdrix):

And the magnitude of the resultant vector F will be,\[\Large |\vec F| \quad=\quad \sqrt{1.519^2+190.57^2}\] \[\Large |\vec F| \quad=\quad 190.576\]

zepdrix (zepdrix):

Ya looks fine :O

zepdrix (zepdrix):

It would be a good idea to make sure you understand these notes though! :) lol

OpenStudy (anonymous):

thank you!, i just wanted to make sure i had it correctly